MICROWAVE SPECTRA OF ISOTOPIC SPECIES OF METHANOL

\underline{KOJIRO TAKAGI}, YUN-BO DUAN and SHOZO TSUNEKAWA, Department of Physics, Toyama University, Toyama 930-8555, Japan.

Microwave spectrum of CH_{3} ^{18}OH has been observed in frequency range between 8 and 200 GHz. Observed lines are those with J ≤ 39, —K—≤8 for both the A and E species in the torsional ground state and some lines, mainly a-type transitions, in the torsional excited states for v_{t}=1-3. Total 568 lines have been fitted to 64 parameters in a reduced Hamiltonian with a root-mean square (RMS) of 0.262 MHz, where the A and E substates are treated simultaneously. Microwave spectrum of CH_{3}OD has been observed in the frequency range between 7 and 200 GHz. These lines have been analyzed by using a reduced Hamiltonian.